# Partial-Label Learning with Conformal Candidate Cleaning

### v1

@@ 1 Introduction
Real-world data is often noisy and ambiguous. In crowd-sourcing, for example, different annotators can assign several conflicting class labels to the same instance. Other examples with ambiguous data include web mining (Guillaumin et al., 2010; Zeng et al., 2013) and audio classification (Briggs et al., 2012). While such datasets can be manually cleaned, sanitizing data is costly, especially for large-scale datasets. Partial-label learning (PLL; Jin and Ghahramani 2002; Lv et al. 2020; Xu et al. 2021; Tian et al. 2024) provides a principled way of dealing with such conflicting data. More specifically, in PLL, instances are annotated with sets of candidate labels of which only one unknown label is the correct class label. PLL permits training a multi-class classifier in this weakly-supervised setting.
@@
  • Experiments. Extensive experiments on artificial and real-world partially labeled data support our claims. An ablation study further demonstrates the usefulness of the proposed strategy. We make our source code and data openly available at github.com/anon1248/pll-with-conformal-candidate-cleaning.
@@
Section 2 establishes our notations and states the partial-label learning problem, Section 3 discusses related work, Section 4 details our contributions, and Section 5 shows our experimental setup and results. All proofs are deferred to Appendix A. Appendix D discusses the method’s runtime complexity and Appendix E lists all hyperparameters used within our experiments in detail.
@@ 4 PLL with Conformal Cleaning
In the following sections, we discuss our method in detail. Section 4.1 elaborates on the notion of conformal validity in the PLL context, Section 4.2 details how to correct for the ambiguity in PLL compared to the supervised setting, Section 4.3 outlines the proposed algorithm, and Section 4.4 discusses the placement of our method with respect to related work.
@@ 4.2 Correcting for Misclassification
Recall that, in the PLL setting, the ground-truth labels $y$ are unavailable during training, which hinders the approximation of (9) needed for the solution of (6). Because a solution to (6) is, however, also desirable in the PLL setting (Theorem 4.1), we make use of existing PLL algorithms to generate pseudo-labels. This strategy iteratively learns a prediction model $f:\operatorname{\mathcal{X}}\to[0,1]^{k}$ that minimizes the empirical risk in (3). We use the trained model $f$ to predict the labels on the validation set $\operatorname{\mathcal{D}}_{v}$, which in turn is used for the calibration of the validity guarantee. Notably, this strategy results in a valid conformal predictor (Theorem 4.4).
@@
Rather than the corrected error level $\alpha^{\prime}_{n}$ as in Theorem 4.4, Remark 4.5 provides a conformal validity guarantee for a fixed error level $\alpha_{2}\in(0,1)$ by using an altered $\alpha^{\prime\prime}$ in the computation of $t_{\alpha^{\prime\prime}}$. If $\alpha^{\prime\prime}\leq 0$, the resulting conformal predictor defaults to $C(x)=\operatorname{\mathcal{Y}}$, for $x\in\operatorname{\mathcal{X}}$, which is valid.
@@ 4.3 Proposed Algorithm
In phase 3 (Lines 11–18), we compute the conformal predictor $C$, which is used to clean the candidate sets. After completing $R_{\operatorname{warmup}}$ warm-up epochs, we start with our pruning procedure. In Line 13, we compute $\alpha_{r}$ for the current epoch $r$. As the constant terms in Theorem 4.4 are unknown, we set $\alpha$ as detailed in the following. Given $(x_{i},s_{i})\in\operatorname{\mathcal{D}}_{t}$, we set $\alpha_{r}$ based on the probability mass that $f$ allocates to non-candidate labels. The more probability mass $f$ puts on non-candidate labels, the more uncertain we are about its predictions and the more we have to correct the error level $\alpha$ to obtain conformal validity. We set $\alpha_{r}=\frac{1}{n^{\prime}}\sum_{i=1}^{n^{\prime}}\sum_{j\notin s_{i}}f_{j}(x_{i})$ with $n^{\prime}=|\operatorname{\mathcal{D}}_{t}|$. Then, we compute the conformal prediction sets $C(x_{i})$ in Line 15 for all training instances $(x_{i},s_{i})\in\operatorname{\mathcal{D}}_{t}$ using the empirical distribution function of the adapted scores on the validation set; this conformal predictor $C$ is valid by Theorem 4.4. We use the conformal sets $C(x_{i})$ to prune the candidate sets $s_{i}$. If $C(x_{i})$ and $s_{i}$ have a nonempty intersection, which is implied with high probability by the conformal validity (Theorem 4.1), we assign $s_{i}\cap C(x_{i})$ to $s_{i}$ in Line 17.
4.4 Placement w.r.t. Related Work
$f_{\hat{y}_{x}}(x)-f_{\hat{o}_{x}}(x)\geq e$ (19)
[…]
5 Experiments
5.1 Algorithms for Comparison
In our experiments, we benchmark six state-of-the-art PLL methods. These are Proden (Lv et al., 2020), Cc (Feng et al., 2020), Valen (Xu et al., 2021), Cavl (Zhang et al., 2022a), Pop (Xu et al., 2023), and CroSel (Tian et al., 2024). For each dataset, we use the same base models across all approaches. For the colored-image datasets, we use a ResNet-9 architecture (He et al., 2016). Else, we use a standard $d$-300-300-300-$k$ MLP (Werbos, 1974). An in-depth overview of all hyperparameters is in Appendix E.
5.2 Experimental Setup
Data.
Using the standard PLL experimentation protocol (Lv et al., 2020; Zhang et al., 2022a; Tian et al., 2024), we perform experiments on real-world PLL datasets and on supervised datasets with artificially added incorrect candidate labels. To report averages and standard deviations, we repeat all experiments five times with different seeds. For the supervised multi-class datasets, we use mnist (LeCun et al., 1999), fmnist (Xiao et al., 2018), kmnist (Clanuwat et al., 2018), cifar10 (Krizhevsky, 2009), cifar100 (Krizhevsky, 2009), and svhn (Netzer et al., 2011). Regarding the real-world PLL datasets, we use bird-song (Briggs et al., 2012), lost (Cour et al., 2011), mir-flickr (Huiskes and Lew, 2008), msrc-v2 (Liu and Dietterich, 2012), soccer (Zeng et al., 2013), and yahoo-news (Guillaumin et al., 2010). An overview of the dataset characteristics is in Appendix E.
Candidate generation.
5.3 Results
@@ Acknowledgements.
This work was supported by the German Research Foundation (DFG) Research Training Group GRK 2153: Energy Status Data — Informatics Methods for its Collection, Analysis and Exploitation and by the KiKIT (The Pilot Program for Core-Informatics at the KIT) of the Helmholtz Association.
@@ Appendix C External Results
Appendix D Runtime Complexity
The main runtime cost of our cleaning method arises from the computation of the conformal sets $C(x_{i})$ in Line 15 of Algorithm 1. Finding the rank of $f_{y}(x_{i})$ within $\operatorname{\mathcal{S}}$ can be done by first sorting $\operatorname{\mathcal{S}}$ and then using a binary search. This requires a total runtime of $\mathcal{O}(Rn\log n)$ as we prune candidate labels in each epoch and, both, the training set $\operatorname{\mathcal{D}}_{t}$ and validation set $\operatorname{\mathcal{D}}_{v}$ have a size of $\mathcal{O}(n)$.
Appendix E Additional Setup
@@
We have implemented all approaches in Python using the Pytorch library. Running all experiments requires approximately three days on a machine with 48 cores and one NVIDIA GeForce RTX 4090. All our source code and data is available at https://github.com/anon1248/pll-with-conformal-candidate-cleaning.

### v2

@@ 1 Introduction
Real-world data is often noisy and ambiguous. In crowdsourcing, for example, different annotators can assign several conflicting class labels to the same instance. Other examples with ambiguous data include web mining (Guillaumin et al., 2010; Zeng et al., 2013) and audio classification (Briggs et al., 2012). While such datasets can be manually cleaned, sanitizing data is costly, especially for large-scale datasets. Partial-label learning (PLL; Jin and Ghahramani 2002; Lv et al. 2020; Xu et al. 2021; Tian et al. 2024) provides a principled way of dealing with such conflicting data. More specifically, in PLL, instances are annotated with sets of candidate labels of which only one unknown label is the correct class label. PLL permits training a multi-class classifier in this weakly-supervised setting.
@@
  • Experiments. Extensive experiments on artificial and real-world partially labeled data support our claims. An ablation study further demonstrates the usefulness of the proposed strategy. We make our source code and data openly available at https://github.com/mathefuchs/pll-with-conformal-candidate-cleaning.
@@
Section 2 establishes our notations and states the partial-label learning problem, Section 3 discusses related work, Section 4 details our contributions, and Section 5 shows our experimental setup and results. All proofs are deferred to Appendix A. Appendix D lists all hyperparameters used within our experiments in detail and Appendix E contains additional experiments.
@@ 4 PLL with Conformal Cleaning
In the following sections, we discuss our method in detail. Section 4.1 elaborates on the notion of conformal validity in the PLL context, Section 4.2 details how to correct for the ambiguity in PLL compared to the supervised setting, Section 4.3 outlines the proposed algorithm, Section 4.4 discusses the method’s runtime complexity, and Section 4.5 discusses the placement of our method with respect to related work.
@@ 4.2 Correcting for Misclassification
Recall that, in the PLL setting, the ground-truth labels $y$ are unavailable during training, which hinders the approximation of (9) needed for the solution of (6). Because a solution to (6) is, however, also desirable in the PLL setting (Theorem 4.1), we make use of existing PLL algorithms to generate pseudo-labels. This strategy iteratively learns a prediction model $f:\operatorname{\mathcal{X}}\to[0,1]^{k}$ that minimizes the empirical risk in (3). We use the trained model $f$ to predict the labels on the validation set $\operatorname{\mathcal{D}}_{v}$, which in turn is used for the calibration of the validity guarantee. Notably, this strategy results in a valid conformal predictor (Theorem 4.4). We note that it remains open to establish the minimality of the resulting conformal sets (analogous to solutions of (6)).
@@
While Remark 4.5 follows from a simple substitution, it explicitly links Theorem 4.4 to the setting usually considered in conformal prediction: One wants to have a conformal predictor that is valid regarding some specified confidence level $\alpha_{2}$, which Remark 4.5 achieves by using an altered $\alpha^{\prime\prime}$ in the computation of $t_{\alpha^{\prime\prime}}$. If $\alpha^{\prime\prime}\leq 0$, the resulting conformal predictor defaults to $C(x)=\operatorname{\mathcal{Y}}$, for $x\in\operatorname{\mathcal{X}}$, which is valid. In contrast, given some confidence level $\alpha$, Theorem 4.4 gives a conformal predictor that is valid with the confidence level $\alpha^{\prime}_{n}\neq\alpha$.
@@ 4.3 Proposed Algorithm
In phase 3 (Lines 11–18), we compute the conformal predictor $C$, which is used to clean the candidate sets. After completing $R_{\operatorname{warmup}}$ warm-up epochs, we start with our pruning procedure. In Line 13, we compute $\alpha_{r}$ for the current epoch $r$. While it is desirable to use the exact value of $\alpha^{\prime\prime}$ in (18) in Line 13 of Algorithm 1, its computation is unfortunately infeasible as the constants $B$ and $\beta$, for which the Bernstein condition (Assumption 4.2) holds, cannot be known unless the true distribution $\operatorname{\mathbb{P}}_{XY}$ is known. As the employed PLL classifiers are consistent, that is, they converge to the Bayes classifier with enough samples, we approximate the estimation error terms in (18) by the probability mass that the PLL classifier allocates on false class labels, that is, class labels that are not part of the candidate sets and hence cannot be the correct label. Given $(x_{i},s_{i})\in\operatorname{\mathcal{D}}_{t}$, we set $\alpha_{r}=\frac{1}{n^{\prime}}\sum_{i=1}^{n^{\prime}}\sum_{j\notin s_{i}}f_{j}(x_{i})$ with $n^{\prime}=|\operatorname{\mathcal{D}}_{t}|$. Then, we compute the conformal prediction sets $C(x_{i})$ in Line 15 for all training instances $(x_{i},s_{i})\in\operatorname{\mathcal{D}}_{t}$ using the empirical distribution function of the adapted scores on the validation set; this conformal predictor $C$ is valid by Theorem 4.4. We use the conformal sets $C(x_{i})$ to prune the candidate sets $s_{i}$. If $C(x_{i})$ and $s_{i}$ have a nonempty intersection, which is implied with high probability by the conformal validity (Theorem 4.1), we assign $s_{i}\cap C(x_{i})$ to $s_{i}$ in Line 17.
4.4 Runtime Complexity
The main runtime cost of our cleaning method arises from the computation of the conformal sets $C(x_{i})$ in Line 15 of Algorithm 1. Finding the rank of $f_{y}(x_{i})$ within $\operatorname{\mathcal{S}}$ can be done by first sorting $\operatorname{\mathcal{S}}$ and then using a binary search. This requires a total runtime of $\mathcal{O}(Rn\log n)$, as we prune candidate labels in each epoch and, both, the training set $\operatorname{\mathcal{D}}_{t}$ and validation set $\operatorname{\mathcal{D}}_{v}$ have a size of $\mathcal{O}(n)$. Note that the runtime of our method is not dependent on the number of feature dimensions $d$ as the considered scores $\operatorname{\mathcal{S}}$ are scalars.
4.5 Placement regarding Related Work
$f_{\hat{y}_{x}}(x)-f_{\hat{o}_{x}}(x)\geq e$ (19)
[…]
5 Experiments
5.1 Algorithms for Comparison
In our experiments, we benchmark six state-of-the-art PLL methods. These are Proden (Lv et al., 2020), Cc (Feng et al., 2020), Valen (Xu et al., 2021), Cavl (Zhang et al., 2022a), Pop (Xu et al., 2023), and CroSel (Tian et al., 2024). For each dataset, we use the same base models across all approaches. For the colored-image datasets, we use a ResNet-9 architecture (He et al., 2016). Else, we use a standard $d$-300-300-300-$k$ MLP (Werbos, 1974). We train all models from scratch. An in-depth overview of all hyperparameters is in Appendix D. Appendix E contains additional experiments, including the use of the pre-trained Blip-2 model (Li et al., 2023) on the vision datasets.
5.2 Experimental Setup
Data.
Using the standard PLL experimentation protocol (Lv et al., 2020; Zhang et al., 2022a; Tian et al., 2024), we perform experiments on real-world PLL datasets and on supervised datasets with artificially added incorrect candidate labels. To report averages and standard deviations, we repeat all experiments five times with different seeds. For the supervised multi-class datasets, we use mnist (LeCun et al., 1999), fmnist (Xiao et al., 2018), kmnist (Clanuwat et al., 2018), cifar10 (Krizhevsky, 2009), cifar100 (Krizhevsky, 2009), and svhn (Netzer et al., 2011). Regarding the real-world PLL datasets, we use bird-song (Briggs et al., 2012), lost (Cour et al., 2011), mir-flickr (Huiskes and Lew, 2008), msrc-v2 (Liu and Dietterich, 2012), soccer (Zeng et al., 2013), and yahoo-news (Guillaumin et al., 2010). An overview of the dataset characteristics is in Appendix D.
Candidate generation.
5.3 Results
@@ Acknowledgements.
This work was supported by the German Research Foundation (DFG) Research Training Group GRK 2153: Energy Status Data — Informatics Methods for its Collection, Analysis and Exploitation and by the pilot program Core-Informatics of the Helmholtz Association (HGF).
@@ Appendix C External Results
Appendix D Additional Setup
@@
We have implemented all approaches in Python using the Pytorch library. Running all experiments requires approximately three days on a machine with 48 cores and one NVIDIA GeForce RTX 4090. All our source code and data is available at github.com/mathefuchs/pll-with-conformal-candidate-cleaning.
Appendix E Additional Experiments
Table 4: Average test-set accuracies ($\pm$ std.) on the real-world datasets. We benchmark our strategy (Conf+) as well as the cleaning method Clsp combined with all existing methods.
Method bird-song lost mir-flickr msrc-v2 soccer yahoo-news
Clsp+Proden 74.61 ($\pm$ 0.84) 61.95 ($\pm$ 2.80) 60.53 ($\pm$ 2.95) 51.74 ($\pm$ 1.77) 31.93 ($\pm$ 29.15) 50.92 ($\pm$ 0.66)
Clsp+Cc 74.37 ($\pm$ 0.91) 60.88 ($\pm$ 3.71) 59.79 ($\pm$ 2.29) 49.64 ($\pm$ 2.06) 53.71 ($\pm$ 0.99) 49.89 ($\pm$ 0.30)
Clsp+Valen 74.95 ($\pm$ 0.27) 59.03 ($\pm$ 2.67) 60.11 ($\pm$ 1.95) 49.92 ($\pm$ 1.80) 53.31 ($\pm$ 0.84) 49.50 ($\pm$ 0.76)
Clsp+Cavl 73.13 ($\pm$ 1.23) 58.76 ($\pm$ 1.75) 59.86 ($\pm$ 2.92) 48.65 ($\pm$ 2.31) 53.48 ($\pm$ 0.76) 49.48 ($\pm$ 0.37)
Clsp+Pop 74.25 ($\pm$ 0.89) 60.18 ($\pm$ 2.48) 59.61 ($\pm$ 1.84) 50.58 ($\pm$ 1.47) 32.08 ($\pm$ 29.29) 50.77 ($\pm$ 0.42)
Clsp+CroSel 76.53 ($\pm$ 1.34) 63.72 ($\pm$ 2.23) 59.75 ($\pm$ 2.79) 51.29 ($\pm$ 1.69) 52.24 ($\pm$ 0.84) 53.53 ($\pm$ 0.93)
Table 5: Average test-set accuracies ($\pm$ std.) on the supervised datasets with added incorrect candidate labels. We benchmark our strategy (Conf) as well as the cleaning method Clsp combined with all existing methods. We use the pre-trained Blip-2 model for all results in this table.
Method mnist fmnist kmnist cifar10 cifar100
Proden 87.21 ($\pm$ 0.83) 71.18 ($\pm$ 2.95) 59.31 ($\pm$ 1.22) 99.07 ($\pm$ 0.05) 90.51 ($\pm$ 0.18)
Clsp+Proden 85.91 ($\pm$ 2.42) 72.11 ($\pm$ 2.81) 62.61 ($\pm$ 1.00) 99.03 ($\pm$ 0.07) 90.31 ($\pm$ 0.13)
Conf+P. (no correction) 91.74 ($\pm$ 0.34) 78.38 ($\pm$ 0.50) 66.88 ($\pm$ 0.76) 99.03 ($\pm$ 0.04) 91.16 ($\pm$ 0.10)
Conf+Proden 91.55 ($\pm$ 0.23) 78.09 ($\pm$ 0.33) 66.43 ($\pm$ 0.38) 99.03 ($\pm$ 0.04) 91.16 ($\pm$ 0.10)
Cc 86.29 ($\pm$ 2.18) 66.19 ($\pm$ 2.77) 58.29 ($\pm$ 0.32) 99.07 ($\pm$ 0.05) 73.43 ($\pm$ 1.40)
Clsp+Cc 85.46 ($\pm$ 1.93) 71.37 ($\pm$ 2.34) 61.37 ($\pm$ 1.09) 99.03 ($\pm$ 0.06) 89.00 ($\pm$ 1.56)
Conf+Cc 85.20 ($\pm$ 4.16) 59.75 ($\pm$ 2.68) 57.07 ($\pm$ 0.66) 99.04 ($\pm$ 0.03) 71.45 ($\pm$ 0.94)
Valen 78.91 ($\pm$ 0.80) 66.53 ($\pm$ 2.65) 58.48 ($\pm$ 0.45) 92.17 ($\pm$ 0.54) 67.24 ($\pm$ 2.49)
Clsp+Valen 84.72 ($\pm$ 3.10) 68.84 ($\pm$ 1.49) 60.76 ($\pm$ 0.76) 98.17 ($\pm$ 0.17) 84.53 ($\pm$ 1.23)
Conf+Valen 74.20 ($\pm$ 21.99) 69.09 ($\pm$ 2.71) 60.95 ($\pm$ 2.59) 42.63 ($\pm$ 19.92) 60.44 ($\pm$ 1.91)
Cavl 71.11 ($\pm$ 3.92) 59.85 ($\pm$ 6.49) 48.15 ($\pm$ 5.07) 41.78 ($\pm$ 21.40) 31.95 ($\pm$ 1.80)
Clsp+Cavl 83.72 ($\pm$ 3.57) 67.38 ($\pm$ 2.59) 62.06 ($\pm$ 2.12) 87.34 ($\pm$ 12.71) 68.02 ($\pm$ 1.96)
Conf+Cavl 71.86 ($\pm$ 4.57) 59.54 ($\pm$ 6.62) 52.14 ($\pm$ 3.89) 29.97 ($\pm$ 15.73) 37.34 ($\pm$ 2.59)
Pop 87.08 ($\pm$ 0.58) 72.30 ($\pm$ 2.63) 60.63 ($\pm$ 1.15) 99.06 ($\pm$ 0.04) 90.50 ($\pm$ 0.21)
Clsp+Pop 85.43 ($\pm$ 2.60) 72.05 ($\pm$ 2.41) 62.49 ($\pm$ 0.90) 99.04 ($\pm$ 0.07) 90.37 ($\pm$ 0.06)
Conf+Pop 91.19 ($\pm$ 0.29) 79.15 ($\pm$ 1.23) 67.37 ($\pm$ 0.28) 99.05 ($\pm$ 0.04) 91.12 ($\pm$ 0.09)
CroSel 91.84 ($\pm$ 0.44) 76.34 ($\pm$ 1.21) 65.55 ($\pm$ 0.81) 99.07 ($\pm$ 0.02) 75.86 ($\pm$ 2.26)
Clsp+CroSel 91.70 ($\pm$ 0.62) 74.42 ($\pm$ 1.02) 67.93 ($\pm$ 1.07) 99.08 ($\pm$ 0.02) 88.80 ($\pm$ 0.85)
Conf+CroSel 91.85 ($\pm$ 0.61) 77.31 ($\pm$ 0.46) 64.73 ($\pm$ 1.52) 99.07 ($\pm$ 0.03) 77.26 ($\pm$ 0.98)
[…]
Comparison vs. all others Wins Ties Losses
Proden 27 37 08
Clsp+Proden 18 27 27
Conf+Proden (no correction) 41 25 06
Conf+Proden 50 20 02
Cc 18 39 15
Clsp+Cc 18 22 32
Conf+Cc 22 30 20
Valen 05 30 37
Clsp+Valen 17 15 40
Conf+Valen 05 24 43
Cavl 05 26 41
Clsp+Cavl 09 19 44
Conf+Cavl 04 26 42
Pop 29 39 04
Clsp+Pop 17 23 32
Conf+Pop 49 22 01
CroSel 30 33 09
Clsp+CroSel 25 15 32
Conf+CroSel 44 22 06
In addition to our cleaning method (Conf), we also benchmark the existing cleaning method Clsp (He et al., 2024) on all datasets in Table 4, 5, and 6, similar to the experiments in Section 5. Instead of training a ResNet-9 base model from scratch as done in Section 5.1, we use the pre-trained Blip-2 model (Li et al., 2023) for the experiments in Table 5 below. We repeat all experiments five times and report means and standard deviations.
We observe that the Clsp models perform well on the image datasets (e.g., cifar100) but poorly on the real-world tabular PLL datasets shown in Table 4. We attribute this to the fact that Clsp relies on the latent representation of large-scale vision models. In contrast, our method Conf gives strong results on, both, real-world and image data. This hypothesis is supported by Table 6: The approaches Conf+Proden, Conf+Pop, and Conf+CroSel that combine the respective approaches with our candidate cleaning strategy win most frequently.